one brother is 5 years older than the other. the sum of thier ages is 27. what are thier ages?
Answer by algebrahouse.com(1659)

x = one brother
x + 5 = other brother

x + x + 5 = 27 {the sum of their ages is 27
2x + 5 = 27 {combined like terms}
2x = 22 {subtracted 5 from each side}
x = 11 {divided each side by 2}
x + 5 = 16 {substituted 5, in for x, into x + 5}

one brother is 11
the other brother is 16

For more help from me, visit: www.algebrahouse.com